# Differentiable and accelerated spherical harmonic and Wigner transforms

### v1

@@ 3 $\times$ GPU Ratio
$L$
Time SO3
Time S2FFT
Speed-Up 1 $\times$ GPU
Speed-Up 3 $\times$ GPU
3 $\times$ GPU Ratio
Time S2FFT
Speed-Up 1 $\times$ GPU
$8$
$3.0\times 10^{-1}$
$3.9\times 10^{0}$
0.08
0.12
1.5
$4.3\times 10^{-2}$
6.98
$16$
$1.1\times 10^{0}$
$7.7\times 10^{0}$
0.14
0.28
2.0
$3.9\times 10^{-2}$
28.7
$32$
$3.4\times 10^{0}$
$1.6\times 10^{1}$
0.21
0.47
2.24
$9.2\times 10^{-2}$
36.3
$64$
$2.9\times 10^{1}$
$3.2\times 10^{1}$
0.93
2.11
2.27
$3.1\times 10^{-1}$
93.9
$128$
$1.3\times 10^{2}$
$6.8\times 10^{1}$
1.96
4.93
2.52
$1.4\times 10^{0}$
94.4
$256$
$1.0\times 10^{3}$
$1.6\times 10^{2}$
6.26
14.9
2.38
$5.9\times 10^{0}$
171
$512$
$6.0\times 10^{3}$
$6.2\times 10^{2}$
9.73
24.9
2.56
$3.0\times 10^{1}$
200
$1024$
$5.8\times 10^{4}$
$4.5\times 10^{3}$
12.9
36.1
2.80
$2048$
$6.4\times 10^{5}$
$3.5\times 10^{4}$
18.4
53.6
2.91
$4096$
$1.0\times 10^{7}$
$2.8\times 10^{5}$
37.1
111
2.99
@@
Fig. 7: Numerical benchmarking results for the Wigner transforms implemented S2FFT compared to the C implementation of SO3. For this analysis we consider MW and MWSS McEwen and Wiaux (2011) samplings of the rotation group. As these samplings afford sampling theorems on the rotation group, their corresponding Wigner transforms are exact. In these comparisons we fix the azimuthal bandlimit to $N=5$, a typical use case corresponding to 9 $\gamma$ rotation angles. S2FFT provides multiple operational modes: a precompute approach with faster compute at $\mathcal{O}(NL^{3})$ memory overhead and an on-the-fly transform with negligible memory overhead (see Sec. 4). Top: Average compute time required for a Wigner transform for each operational modality. Middle: Relative acceleration of a given modality with respect to their SO3 counterpart. Bottom: Averaged round-trip numerical error. In each case we recover machine precision. We also empirically observe at most $\mathcal{O}(L)$ error scaling (in line with previous work McEwen et al. (2015)). The silver and gold stars represent the performance one may expect to realise should one distribute our on-the-fly algorithms across 100 and 1000 GPUs respectively.
@@ 7 Conclusions
We have developed differentiable and accelerated algorithms to compute generalised Fourier transforms on the sphere and rotation group and their gradients. We present a recursive algorithm for the calculation of Wigner $d$-functions that is both stable to high harmonic degrees and extremely parallelisable. Leveraging this recursion we develop fast and exact spin spherical harmonic and Wigner transforms that are well-suited for the high throughput computing provided by modern hardware accelerators. We develop a hybrid differentiation approach so that gradients can be computed efficiently, avoiding the memory overhead of full automatic differentiation while also avoiding the complexities of full manual differentiation. Both forward- and reverse-mode differentiation is supported.
@@
S2FFT is comprehensively benchmarked up to bandlimit $L=8192$ (higher bandlimits are certainly feasible but have not been considered to date due to the computational burden of further benchmarking). The transforms and gradient computations implemented are validated, where for sampling theorems that admit exact spherical transforms the error of a round-trip transform is observed to be of order of machine precision. In addition we benchmark the accelerated computation time of S2FFT against alternative C codes implementing spherical harmonic McEwen and Wiaux (2011) and Wigner McEwen et al. (2015) transforms. Our transforms provide up to a 234-fold and 400-fold acceleration, respectively, for the on-the-fly and precompute transforms. Moreover, our spin spherical harmonic transforms are engineered to distribute compute across many accelerators. For high bandlimits we achieve very close to optimal linear scaling with increasing number of GPUs due to the highly parallelised and balanced nature of our algorithms, i.e. when considering $3$ GPUs we realise a 3×3$\times$3 × acceleration compared to running on a single GPU. Provided access to sufficiently many GPUs our transforms thus exhibit an unprecedented effective linear time complexity.
With researchers becoming increasingly interested in differentiable programming for scientific applications, we hope these foundational tools will be of great use in coming years. As just one example of potential applications, for global spherical modelling or analysis of the Earth’s climate or geophysical properties through data-driven or hybrid methods, sub-kilometre resolution and below is now feasible.

### v2

@@ 6.3.2 Computational speed
$L$ Time SO3 Time S2FFT Speed-Up 1 $\times$ GPU Speed-Up 3 $\times$ GPU 3 $\times$ GPU Ratio Time S2FFT Speed-Up 1 $\times$ GPU
$8$ $3.0\times 10^{-1}$ $3.9\times 10^{0}$ 0.08 0.12 1.5 $4.3\times 10^{-2}$ 6.98
$16$ $1.1\times 10^{0}$ $7.7\times 10^{0}$ 0.14 0.28 2.0 $3.9\times 10^{-2}$ 28.7
$32$ $3.4\times 10^{0}$ $1.6\times 10^{1}$ 0.21 0.47 2.24 $9.2\times 10^{-2}$ 36.3
$64$ $2.9\times 10^{1}$ $3.2\times 10^{1}$ 0.93 2.11 2.27 $3.1\times 10^{-1}$ 93.9
$128$ $1.3\times 10^{2}$ $6.8\times 10^{1}$ 1.96 4.93 2.52 $1.4\times 10^{0}$ 94.4
$256$ $1.0\times 10^{3}$ $1.6\times 10^{2}$ 6.26 14.9 2.38 $5.9\times 10^{0}$ 171
$512$ $6.0\times 10^{3}$ $6.2\times 10^{2}$ 9.73 24.9 2.56 $3.0\times 10^{1}$ 200
$1024$ $5.8\times 10^{4}$ $4.5\times 10^{3}$ 12.9 36.1 2.80
$2048$ $6.4\times 10^{5}$ $3.5\times 10^{4}$ 18.4 53.6 2.91
$4096$ $1.0\times 10^{7}$ $2.8\times 10^{5}$ 37.1 111 2.99
@@
Fig. 7: Numerical benchmarking results for the Wigner transforms implemented S2FFT compared to the C implementation of SO3. For this analysis we consider MW and MWSS [41] samplings of the rotation group. As these samplings afford sampling theorems on the rotation group, their corresponding Wigner transforms are exact. In these comparisons we fix the azimuthal bandlimit to $N=5$, a typical use case corresponding to 9 $\gamma$ rotation angles. S2FFT provides multiple operational modes: a precompute approach with faster compute at $\mathcal{O}(NL^{3})$ memory overhead and an on-the-fly transform with negligible memory overhead (see Sec. 4). Top: Average compute time required for a Wigner transform for each operational modality. Middle: Relative acceleration of a given modality with respect to their SO3 counterpart. Bottom: Averaged round-trip numerical error. In each case we recover machine precision. We also empirically observe at most $\mathcal{O}(L)$ error scaling (in line with previous work [42]). The silver and gold stars represent the performance one may expect to realise should one distribute our on-the-fly algorithms across 100 and 1000 GPUs respectively.
@@ 7 Conclusions
We have developed differentiable and accelerated algorithms to compute generalised Fourier transforms on the sphere and rotation group and their gradients. We present a recursive algorithm for the calculation of Wigner $d$-functions that is both stable to high harmonic degrees and extremely parallelisable. This level of parallelisation is achieved as we recurse along harmonic order $m$ independently for all other indices, which is stable due to a custom on-the-fly normalisation scheme. Leveraging this recursion we develop fast and exact spin spherical harmonic and Wigner transforms that are well-suited for the high throughput computing provided by modern hardware accelerators. We develop a hybrid differentiation approach so that gradients can be computed efficiently, avoiding the memory overhead of full automatic differentiation while also avoiding the complexities of full manual differentiation. Both forward- and reverse-mode differentiation is supported.
@@
S2FFT is comprehensively benchmarked up to bandlimit $L=8192$ (higher bandlimits are certainly feasible but have not been considered to date due to the computational burden of further benchmarking). The transforms and gradient computations implemented are validated, where for sampling theorems that admit exact spherical transforms the error of a round-trip transform is observed to be of order of machine precision. In addition we benchmark the accelerated computation time of S2FFT against alternative C codes implementing spherical harmonic [41] and Wigner [42] transforms. Our transforms provide up to a 234-fold and 400-fold acceleration, respectively, for the on-the-fly and precompute transforms. Moreover, our spin spherical harmonic transforms are engineered to distribute compute across many accelerators. For high bandlimits we achieve very close to optimal linear scaling with increasing number of GPUs due to the highly parallelised and balanced nature of our algorithms, i.e. when considering $3$ GPUs we realise a 3×3$\times$3 × acceleration compared to running on a single GPU. Provided access to sufficiently many GPUs our transforms thus exhibit an unprecedented effective linear time complexity.
With researchers becoming increasingly interested in differentiable programming for scientific applications, we hope these foundational tools will be of great use in coming years. The differentiability afforded by our transforms means that, for example, scattering transforms on the sphere [33], which rely on spherical harmonic, Wigner, and spherical wavelet transforms, can be extended to differentiable transforms, which then opens up their use as statistical generative models, as has been considered previously in the planar setting [e.g. 76]. Indeed, we have already leveraged the differentiable spherical harmonic and Wigner transforms developed in the current article to realise differentiable and accelerated wavelet [77] and scattering [78] transforms on the sphere, and corresponding generative models. Furthermore, the acceleration afforded by our transforms means that, for example, sub-kilometre resolution and below is now feasible for global spherical modelling or analysis of the Earth’s climate or geophysical properties through data-driven or hybrid methods. These are just a small number of potential uses of the differentiable and accelerated spherical transforms developed in the current article; we hope they will find many more uses in the sciences and beyond.